### v2

@@ 4.3 Summary
Our survey was conducted to complement the SLR findings with real-world insights from industry experts. The survey involved 32 practitioners, predominantly senior-level professionals with over ten years of experience in agile software development. The survey confirmed many of the challenges identified in the SLR regarding data management aspects (RQ1). The results show that managing data integration processes, capturing diverse data, automation data collection challenges, ensuring data accuracy and consistency, completeness of data, complex data sets, and real-time analysis requirements are the most notable challenges faced by the participants. The participants adopted a variety of solutions to address different data management challenges, and those are mainly in line with the solutions reported in the SLR (RQ2). However, the survey results highlighted the need for training programs that focus on data management skills, an aspect that was not given much emphasis in the previous studies. The findings of the survey show a greater need for real-time analytics and automation than indicated in the SLR. Detailed descriptions and examples of each challenge and solution are provided in the Appendix.
@@ 5.1 Comparison of the SLR and the Survey Findings
The survey results also reflect data management challenges identified in the SLR. Most participants reported challenges with managing data integration processes, emphasizing the need for effective management of diverse data sources, which directly aligns with the SLR’s findings on the complexities of managing data integration across various systems. Similarly, most survey respondents highlighted challenges with ensuring data accuracy, consistency and completeness, which supports the SLR’s emphasis on the importance of data quality for effective decision-making.
The survey offered additional practical insights into data collection and analysis challenges. Most practitioners reported challenges regarding capturing diverse data during data collection and highlighted difficulties with automating data collection processes. These findings align with the SLR’s emphasis on the need for comprehensive and standardized data collection methods. Regarding data analysis, most participants reported difficulties with analyzing complex data sets and the need for real-time analysis. This underscores the importance of advanced analytical tools and techniques, including communication-centric approaches and agile workflow adaptations, as solutions for complex data analysis. The survey also showed a notable need for training programs that enhance data management skills within agile teams. This aspect was not given much emphasis in the studies in the SLR. Implementing real-time data analytics proved challenging for practitioners, who noted the need for better tools and training. The survey findings highlight a greater need for automation and real-time analytics than those mentioned in the SLR.
@@ 5.2 Implications
Requirements Gathering: Inaccurate or incomplete data during requirements gathering can lead to poorly defined project goals and user stories, negatively impacting the entire development process and potentially leading to unmet stakeholder expectations. Product owners may struggle to define clear project goals. To mitigate this, they should ensure effective communication with stakeholders and use comprehensive data collection methods. Business analysts may struggle to gather and document project requirements accurately. They should employ structured data-gathering techniques, use appropriate tools, and validate data with multiple sources. The project team may receive poorly defined requirements, leading to potential rework. To mitigate this, the team should collaborate closely with all stakeholders, thoroughly review requirements before implementation, and provide feedback to improve data collection methods.
@@
Retrospectives: Retrospectives aim to reflect on the past iteration to understand what went well, what went wrong, and what could be improved dharmapal2016big ; matthies2020playing . Data management challenges (e.g., challenges in managing the integration process, such as integrating development project data into retrospective agendas) can hinder the agile development team’s ability to analyze project performance data, derive actionable insights, and implement improvements. They should use data integration tools, ensure data accuracy and completeness, and facilitate collaborative analysis sessions. Moreover, scrum masters may struggle to guide the team in reflecting on past iterations and deriving actionable insights. They should use structured retrospective techniques, ensure the availability of accurate data, and guide the team in data analysis. Product owners may face challenges in understanding the overall project performance and making informed decisions. They should use comprehensive data analytics tools, ensure regular data reviews, and involve stakeholders in the retrospective process.
@@ 5.3 Recommendations for Practitioners
We recommend developing comprehensive data management policies to address the data management challenges in support of agile methodologies used by practitioners. The data management policies require a collaborative effort among multiple stakeholders. Data governance team can lead the policy creation, ensuring alignment with regulatory requirements and organizational goals. Our SLR and survey show that including roles in the agile development team, such as legal advisors, can provide guidance on legal obligations and help the agile development team adhere to compliance with data privacy regulations barbala2023data . The challenges that pertain to large volumes of data (e.g., ensuring data quality of large datasets) can handled by machine learning developers, who also help ensure that machine learning models are trained on accurate and representative datasets dautov2022towards . Agile coaches assist teams in implementing data-driven decision-making policies, which help effectively utilise data to guide development processes batarseh2018predicting . To develop comprehensive data management policies, we recommend that agile teams review the data management challenges and pick suitable solutions that suit their project and agile environment settings.
@@ 6.1 Internal Validity
To ensure a quality selection of articles, we conducted thorough quality checks to ensure each selected study’s rigour, integrity, and relevance. Two authors evaluated each study using a predetermined quality assessment framework (see Section 3.1.5), resolving any disagreements through discussion kitchenham2007guidelines .
@@
In terms of the practitioner survey, we iteratively designed the survey questions to ensure we designed a relevant and reliable questionnaire. This was done collaboratively between the researchers involved. We then piloted the questionnaire with three practitioners (with experience in data management) to ensure that the questions were relevant and understandable. This was done to ensure the survey was relevant and valid for the target audience. Additionally, we followed a detailed data analysis procedure with verification through meetings to minimize bias. The questionnaire development incorporated feedback from the pilot survey we ran with two agile practitioners with experience in dealing with data.
@@ 6.2 External Validity
In the survey part of the study, we targeted a diverse group of practitioners from various industries and levels of experience in agile software development to ensure diversity and representation of the different roles. This diversity can help enhance our findings’ generalizability to a broad range of agile software development environments. The survey participants were from eight countries working on various agile development projects, which helps to increase the generalizability of the findings.
@@ 6.3 Construct Validity
To reduce data extraction basis for both the SLR and the survey, we standardized data extraction forms to ensure uniformity. We refined our data extraction methods after several pilot tests. Two authors reviewed the extracted data process, resolving discrepancies through consensus. Each article that was extracted was then discussed between the researchers. We maintained a detailed description of the extraction process along with a replication package to ensure transparency and allow other researchers to verify our methods and findings fawzy_2024_10597818 .
The survey questionnaire was designed to accurately measure the intended constructs kitchenham2008personal ; linaaker2015guidelines ; robson2002real . We considered convergent and divergent validity and integrated these into our survey design. We achieved this through the execution of a pilot survey, which helped to ensure the construct and content validity.
@@ 7 Conclusions
In this study, we investigated data management challenges in agile software development and the proposed solutions through an SLR and a complementary survey with practitioners. The SLR identified 45 studies, categorizing various aspects of data management such as data integration, quality, collection, and analysis. We then followed this analysis by surveying 32 practitioners with extensive experience with agile practices. This survey confirmed the main challenges we identified in our SLR and uncovered additional practical challenges that practitioners have reported. Our findings (from both the SLR and survey) reveal that managing data integration processes, capturing diverse data, automation data collection, ensuring data accuracy and consistency, completeness of data, complexity of data, and real-time analysis requirements are the most notable data management challenges faced by practitioners in agile development. Solutions, such as the adoption of ontology-based approaches and the development of cloud-based platforms and automated tools, have shown promising results in mitigating these data management challenges. Furthermore, these data management challenges impact agile process activities and project roles, leading to poorly defined project goals and user stories. Product owners, business analysts and scrum masters should use effective data management techniques to improve decision-making and overall project success.
Future studies should investigate advanced data integration techniques to harmonize and interoperate diverse and heterogeneous data sources. A promising research direction is the use of ontology-based, architecture-centric approaches and the adoption of AI techniques to resolve data management challenges. Furthermore, future studies should investigate systematic and automated data collection methodologies to enhance data accuracy and completeness throughout the software development life-cycle. This includes advanced automated quality assurance techniques, which combine real-time quality metrics and continuous monitoring tools. Such techniques and tools can help manage large volumes of structured and unstructured data and significantly improve agile teams’ decision-making abilities. Future research should also rigorously evaluate the proposed solutions from both the SLR and practitioners’ surveys to assess their effectiveness and applicability in real-world agile environments.
@@ Acknowledgments
We would like to thank all practitioners who participated in our survey for their valuable time and input.

### v4

@@ 4.4 Summary
Our survey was conducted to complement the SLR findings with real-world insights from industry agile practitioners who were significantly involved in data management activities. The survey involved 32 practitioners, predominantly senior-level professionals with over ten years of experience in agile software development. The survey confirmed many of the challenges identified in the SLR regarding data management aspects (RQ1). The results show that managing data integration processes, capturing diverse data, automation data collection challenges, ensuring data accuracy and consistency, completeness of data, complex data sets, and real-time analysis requirements are the most notable challenges faced by the participants. The participants adopted various solutions to address different data management challenges, which are mainly in line with the solutions reported in the SLR (RQ2). However, the survey results highlighted the need for training programs focusing on data management skills, which was not emphasized much in the previous studies. The survey findings show a greater need for real-time analytics and automation than indicated in the SLR. Detailed descriptions and examples of each challenge and solution are provided in the Appendix.
@@ 5.1 Comparison of the SLR and the Survey Findings
The survey results also reflect data management challenges identified in the SLR. Most participants reported challenges with managing data integration processes, emphasizing the need for effective management of diverse data sources, which directly aligns with the SLR’s findings on the complexities of managing data integration across various systems. Similarly, most survey respondents highlighted challenges with ensuring data accuracy, consistency, and completeness, which supports the SLR’s emphasis on the importance of data quality for effective decision-making.
The survey offered additional practical insights into data collection and analysis challenges. Most practitioners reported challenges regarding capturing diverse data during data collection and highlighted difficulties with automating data collection processes. These findings align with the SLR’s emphasis on the need for comprehensive and standardized data collection methods. Regarding data analysis, most participants reported difficulties analyzing complex data sets and needing real-time analysis. This underscores the importance of advanced analytical tools and techniques, including communication-centric approaches and agile workflow adaptations, as solutions for complex data analysis. The survey also showed a notable need for training programs that enhance data management skills within agile teams. This aspect was not given much emphasis in the studies in the SLR. Implementing real-time data analytics proved challenging for practitioners, who noted the need for better tools and training. The survey findings highlight a greater need for automation and real-time analytics than those mentioned in the SLR.
@@ 5.2 Implications
Requirements Gathering: Inaccurate or incomplete data during requirements gathering can lead to poorly defined project goals and user stories, negatively impacting the entire development process and potentially leading to unmet stakeholder expectations. Product owners may struggle to define clear project goals. To mitigate this, they should ensure effective stakeholder communication and use comprehensive data collection methods. Business analysts may struggle to gather and document project requirements accurately. They should employ structured data-gathering techniques, use appropriate tools, and validate data with multiple sources. The project team may receive poorly defined requirements, leading to potential rework. To mitigate this, the team should collaborate closely with all stakeholders, thoroughly review requirements before implementation, and provide feedback to improve data collection methods.
@@
Retrospectives: Retrospectives aims to reflect on the past iteration to understand what went well, what went wrong, and what could be improved dharmapal2016big ; matthies2020playing . Data management challenges (e.g., challenges in managing the integration process, such as integrating development project data into retrospective agendas) can hinder the agile development team’s ability to analyze project performance data, derive actionable insights, and implement improvements. They should use data integration tools, ensure data accuracy and completeness, and facilitate collaborative analysis sessions. Moreover, scrum masters may struggle to guide the team in reflecting on past iterations and deriving actionable insights. They should use structured retrospective techniques, ensure the availability of accurate data, and guide the team in data analysis. Product owners may face challenges in understanding the overall project performance and making informed decisions. They should use comprehensive data analytics tools, ensure regular data reviews, and involve stakeholders in the retrospective process.
@@ 5.3 Recommendations for Practitioners
We recommend developing comprehensive data management policies to address the data management challenges in support of agile methodologies used by practitioners. The data management policies require a collaborative effort among multiple stakeholders. Data governance team can lead the policy creation, ensuring alignment with regulatory requirements and organizational goals. Our SLR and survey show that including roles in the agile development team, such as legal advisors, can provide guidance on legal obligations and help the agile development team adhere to compliance with data privacy regulations barbala2023data . The challenges that pertain to large volumes of data (e.g., ensuring data quality of large datasets) can handled by machine learning developers, who also help ensure that machine learning models are trained on accurate and representative datasets dautov2022towards . Agile coaches assist teams in implementing data-driven decision-making policies, which help effectively utilize data to guide development processes batarseh2018predicting . To develop comprehensive data management policies, we recommend that agile teams review the data management challenges and pick suitable solutions that suit their project and agile environment settings.
@@ 6.1 Internal Validity
To ensure a quality selection of articles, we conducted thorough quality checks to ensure each selected study’s rigor, integrity, and relevance. Two authors evaluated each study using a predetermined quality assessment framework (see Section 3.1.5), resolving any disagreements through discussion kitchenham2007guidelines .
@@
While conducting the SLR, we did not employ a literature snowballing process to screen for related articles cited in the SLR studies. This may have led to potentially missing relevant studies not being captured in our automated search. Although we initially reviewed references from the included studies, the additional actual studies identified were insufficient to justify their inclusion in our SLR. We adopted an informal snowballing approach (by checking the references of a subset of the nine most frequently appearing studies in our search) to cross-verify the results and ensure no significant studies were omitted. This approach involved examining references to enhance the robustness of our dataset.
In terms of the practitioner survey, we iteratively designed the survey questions to ensure we designed a relevant and reliable questionnaire. This was done collaboratively between the researchers involved. We then piloted the questionnaire with three practitioners (with experience in data management) to ensure that the questions were relevant and understandable. This was done to ensure the survey was relevant and valid for the target audience. Additionally, we followed a detailed data analysis procedure with verification through meetings to minimize bias. The questionnaire development incorporated feedback from the pilot survey we ran with two agile practitioners who had experience dealing with data.
@@ 6.2 External Validity
In the survey part of the study, we targeted a diverse group of practitioners from various industries and levels of experience in agile software development to ensure diversity and representation of the different roles. This diversity can help enhance our findings’ generalizability to a broad range of agile software development environments.
The survey participants represented eight countries and worked on various agile development projects. However, we acknowledge the limitations imposed by the small sample size, the overrepresentation of participants from certain countries (e.g., New Zealand and Saudi Arabia), the limited representation from others (e.g., only one respondent each from France, Germany, and the USA), and the predominance of certain roles (e.g., developers and product owners) in our survey sample, which can restrict the generalizability of the findings. Still, we aimed to ensure that we have a diverse range of industry sectors and seniority levels represented in our sample.
@@ 6.3 Construct Validity
To reduce the data extraction basis for both the SLR and the survey, we standardized data extraction forms to ensure uniformity. We refined our data extraction methods after several pilot tests. Two authors reviewed the extracted data process, resolving discrepancies through consensus. Each article that was extracted was then discussed between the researchers. We maintained a detailed description of the extraction process along with a replication package to ensure transparency and allow other researchers to verify our methods and findings fawzy_2024_10597818 . The survey questionnaire was designed to accurately measure the intended constructs kitchenham2008personal ; linaaker2015guidelines ; robson2002real . We considered convergent and divergent validity and integrated these into our survey design. We achieved this through the execution of a pilot survey, which helped to ensure the construct and content validity.
@@ 7 Conclusions
In this study, we investigated data management challenges in agile software development and the proposed solutions through an SLR and a complementary survey with practitioners. The SLR identified 45 studies, categorizing various aspects of data management such as data integration, quality, collection, and analysis. We then followed this analysis by surveying 32 practitioners with extensive experience with agile practices. This survey confirmed the main challenges we identified in our SLR and uncovered additional practical challenges that practitioners have reported. Our findings (from both the SLR and survey) reveal that managing data integration processes, capturing diverse data, automation data collection, ensuring data accuracy and consistency, completeness of data, complexity of data, and real-time analysis requirements are the most notable data management challenges faced by practitioners in agile development. Solutions, such as adopting ontology-based approaches and developing cloud-based platforms and automated tools, have shown promising results in mitigating these data management challenges. Furthermore, these data management challenges impact agile process activities and project roles, leading to poorly defined project goals and user stories. Product owners, business analysts, and scrum masters should use effective data management techniques to improve decision-making and overall project success.
Future studies should investigate advanced data integration techniques to harmonize and interoperate diverse and heterogeneous data sources. A promising research direction is using ontology-based, architecture-centric approaches and adopting AI techniques to resolve data management challenges. Furthermore, future studies should investigate systematic and automated data collection methodologies to enhance data accuracy and completeness throughout the software development life-cycle. This includes advanced automated quality assurance techniques combining real-time quality metrics and continuous monitoring tools. Such techniques and tools can help manage large volumes of structured and unstructured data and significantly improve agile teams’ decision-making abilities. Future research should also rigorously evaluate the proposed solutions from both the SLR and practitioners’ surveys to assess their effectiveness and applicability in real-world agile environments.
Conflicts of Interest
The authors declared that they have no conflict of interest in the submission of this manuscript.
Data Availability Statement
A dataset of both the SLR and the practitioner survey is available online in our replication package fawzy_2024_10597818 .
@@ Acknowledgments
We would like to thank all practitioners who participated in our survey for their valuable time and input. Peng Liang is partially funded by NSFC under Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.